Rental costs for office space have been going up at 9.9% per year compounded annually for the past 5 years. If the office rent is now $29 per square foot per month, what were the innitial rates 5 years ago?
The rates 5 years ago were about $______ per square foot.
(Round to two decimal places as needed.)
Answer #1

Rate 5 year ago can be calculated by solving the below equation:

Rents years ago X (1+r) = Rentnou Rent5 years ago x (1 + 0.099) = 29 29 Rents years ago = (1 + 0.099) Rent5 years ago = 18.1

So the rent was 18.10 per month per foot 5 years ago